OpenStudy (anonymous):

solve cos3A+sinA-cosA=0,for 0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

really? i think you get a cubic equation if you try to solve this.

OpenStudy (anonymous):

i read it as \[\cos(3A)\] but perhaps it is \[\cos^3(A)\] for the first term

OpenStudy (anonymous):

read it as cos(3A)

OpenStudy (anonymous):

then you are in trouble

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

well you have to use a triple angle formula for cosine. then try to rewrite it in term of only one trig function, but i am fairly sure you will get a cubic, and solving will not be easy

OpenStudy (anonymous):

haha thx anywayz but i believe there is a simpler way will get back to u i i got it

OpenStudy (anonymous):

i'd love to see it. pi works. 0 works. if you just want an answer and see how ugly it is try this http://www.wolframalpha.com/input/?i=cos%283A%29%2Bsin%28A%29-cos%28A%29%3D0

OpenStudy (anonymous):

haha thx

OpenStudy (anonymous):

let me know the "easy" way to get the weird looking ones

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey if i think if we change cos3A to cos(2A+A) it might be easier to solve since we can use trigo addition formular

OpenStudy (anonymous):

well cos3A - cos A = -2sin2AsinA so we got -2sin2AsinA + sinA = 0

OpenStudy (anonymous):

how can we proceed now can we divide by sinA?? giving -2sin2A + 1 = 0 so sin2A = 0.5 2A = 30, 150, 390, 510 then A =15, 75, 195, 255 but 0 and 360 are roots also

OpenStudy (anonymous):

is uppose you lose some roots by divivding by sinA

OpenStudy (anonymous):

they are angles heard of the ASTC method??

OpenStudy (anonymous):

how did u get-2sin2Asin2A ??

OpenStudy (anonymous):

ASTC no

OpenStudy (anonymous):

by the 'sums and differences' formulae there are four

OpenStudy (anonymous):

trigo addition formular??

OpenStudy (anonymous):

the one I used is CosA - cos B = -2 sin(A+B)/2 * sin(A-B/2

OpenStudy (anonymous):

read as 'cos - cos = - 2sin semi sum sin semi difference'

OpenStudy (anonymous):

factor formular i see i see

OpenStudy (anonymous):

i'll check some of the results

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

soc45 + sin15 - cos 15 = 0 that was is right anyway

OpenStudy (anonymous):

solved A=0,90,180

OpenStudy (anonymous):

those are the answered given?

OpenStudy (anonymous):

urs correct mine wrong had a careless mistake

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

I had to use one of those formulae a liitle while ago otherwise I wouldn have remembered them They can be useful!! OK u welcome. Pls tick the good answer box to close the question

OpenStudy (anonymous):

haha done thx once again

OpenStudy (anonymous):

nice work jimmyrep! i got \[\frac{\pi}{12}\] by guessing, but did not get the others